OpenStudy (anonymous):

In Karly's history class, 3/10 of the students are making projects for their final project and 1/2 are writing essays. What fraction of the students are making either a project or writing essays?

OpenStudy (anonymous):

If we presume that in Karlys class, students are doing either final project or writing essays and not doing anything else :P, then we have 3/10 making projects and 1/2 writing essays, so remaining people would be 1 - 3/10 -1/2

OpenStudy (anonymous):

if we can make the denominator same, that is 10, it would be easy for us to do calculation . 1 we can divide and mutliply with 10 to it make 10/10 1/2 we can multiply and divide with 5 to make it 5/10 so 10/10 - 3/10 -5/10 = (10 - 3 - 5)/10 = (10 -8)/10 = 2/10 = 1/5